# An operator learning perspective on parameter-to-observable maps

### v1

@@ 1.4. Outline
The remainder of this article is organized as follows. We define the architecture of FNMs as a slight adjustment of FNOs in Section 2 and confirm that FNMs retain desirable properties of FNOs such as universal approximation in Section 3. In Section 4, we analyze end-to-end and full-field learning of linear functionals to establish a theoretical foundation that underlies the data volume requirements of the two approaches. Going beyond the theory, Section 5 provides numerical experiments that compare end-to-end and full-field learning with FNMs with both finite- and infinite-dimensional input space representations for predicting QoIs in several nonlinear PDE problems. Concluding remarks are given in Section 6. Appendix A contains additional theorems related to Section 4. All proofs are provided in Appendices B and C.
2. Neural mappings for finite-dimensional inputs or outputs
@@ 2.1. A review of neural operators
Let $\mathcal{U}=\mathcal{U}(\mathcal{D};\mathbb{R}^{d_{u}})$ and $\mathcal{Y}=\mathcal{Y}(\mathcal{D};\mathbb{R}^{d_{y}})$ be Banach spaces over Euclidean domain $\mathcal{D}\subset\mathbb{R}^{d}$. Finite-dimensional fully-connected neural networks are repeated compositions of affine mappings alternating with pointwise nonlinearities. To extend this framework to the infinite-dimensional function space setting, depth $T$ neural operators from $\mathcal{U}$ to $\mathcal{Y}$ take the form
@@
where $\mathcal{S}$ is a pointwise-defined local lifting operator, $\mathcal{Q}$ is a pointwise-defined local projection operator, and for each $t\in[T]$, the layer $\mathscr{L}_{t}$ is a nonlinear map between function spaces that is the composition of a local (and usually nonlinear) operator with a nonlocal affine kernel integral operator [31].
The Fourier Neural Operator (FNO) is a specific instance of the class of neural operators (1) where, for $t\in[T]$, the form of the layer $\mathscr{L}_{t}\colon\{v\colon\mathbb{T}^{d}\to\mathbb{R}^{d_{t-1}}\}\to\{v\colon\mathbb{T}^{d}\to\mathbb{R}^{d_{t}}\}$ is given by
$v\mapsto\mathscr{L}_{t}(v)=\Bigl{\{}\sigma_{t}\bigl{(}W_{t}v(x)+(\mathcal{K}_{t}v)(x)+b_{t}(x)\bigr{)}\Bigr{\}}_{x\in\mathbb{T}^{d}}\,.$ (2)
@@
In the preceding display, the $\psi_{k}=e^{2\pi i\langle k,{\,\cdot\,}\rangle_{\mathbb{R}^{d}}}$ are the complex Fourier basis elements of $L^{2}(\mathbb{T}^{d};\mathbb{C})$ and $P^{(k)}_{t}\in\mathbb{C}^{d_{{t}}\times d_{t-1}}$ are the learnable parameters of the integral operator $\mathcal{K}_{t}$ for each $k\in\mathbb{Z}^{d}$. The functions $\sigma_{t}\colon\mathbb{R}\to\mathbb{R}$ are nonlinear activations that act pointwise when applied to vectors. Additional details of more general versions and computational implementations of the FNO may be found in [30, 31, 39].
Though the internal FNO layers $\{\mathscr{L}_{t}\}$ in (2) and (3) are defined on the periodic domain $\mathbb{T}^{d}$, it is possible to apply the FNO to other domains $\mathcal{D}\subset\mathbb{R}^{d}$. To do this, introduce an operator :$\mathcal{E}\colon\{h\colon\mathcal{D}\to\mathbb{R}^{d_{0}}\}\to\{h\colon\mathbb{T}^{d}\to\mathbb{R}^{d_{0}}\}$ that maps to functions on $\mathbb{T}^{d}$ and replace $\mathcal{S}$ in (1) with $\mathcal{E}\circ\mathcal{S}$. Similarly, let :$\mathcal{R}\colon\{h\colon\mathbb{T}^{d}\to\mathbb{R}^{d_{T}}\}\to\{h\colon\mathcal{D}\to\mathbb{R}^{d_{T}}\}$ be an operator that maps back to functions on the desired domain $\mathcal{D}$ and replace $\mathcal{Q}$ in (1) with $\mathcal{Q}\circ\mathcal{R}$. These modifications to the lifting and projecting components yield the final FNO architecture
@@ 2.2. The neural mappings framework
The neural operator architecture described in Section 2.1 only accepts inputs, outputs, and intermediate states that are elements of function spaces. Finite-dimensional inputs, outputs, and states are not directly compatible with neural operators. We propose neural mappings, which lift this restriction through two fundamental building blocks. The first, linear functional layers, map from function space to finite dimensions. The second, linear decoder layers, map from finite dimensions to function space. We combine these two building blocks with standard iterative neural operator layers to form several classes of nonlinear and function space consistent architectures.
Instating the neural operator notation from Section 2.1, we define a linear functional layer $\mathscr{G}\colon\{h\colon\mathcal{D}\to\mathbb{R}^{d_{T-1}}\}\to\mathbb{R}^{d_{T}}$ and a linear decoder layer $\mathscr{D}\colon\mathbb{R}^{d_{0}}\to\{h\colon\mathcal{D}\to\mathbb{R}^{d_{1}}\}$ to be maps of the form
@@
where we recall that $\{\psi_{k}\}$ is the Fourier basis of $L^{2}(\mathbb{T}^{d};\mathbb{C})$. In (6), the entries of the matrices $\{P^{(k)}\}\subset\mathbb{C}^{d_{T}\times d_{T-1}}$ correspond to the Fourier coefficients of the function $\kappa$ in (5). Similar calculations show that, on the torus $\mathbb{T}^{d}$, $\mathscr{D}$ takes the form
@@
We are now able to define the general FNMs architecture.
@@ 3. Universal approximation theory for Fourier Neural Mappings
In this section, we establish universal approximation theorems for FNMs; this is a confirmation that the architectures maintain this desirable property of neural operators. The results are stated for the cases of the F2V and V2F architectures; the case of V2V trivially follows. Similar results also hold for general neural mappings by invoking the appropriate universal approximation theorems for general neural operators from [30, Section 9.3] and for the topology induced by Lebesgue-Bochner norms, i.e., average error with respect to a probability measure supported on the input space. For more details regarding these extensions, see [31, Theorems 11–14, Section 9.3, pp. 55–57] and [30, pp. 12–14 and Theorem 18]. Our proofs, which are collected in Appendix B, use arguments based on constant functions that are similar to those used to prove universal approximation theorems at the level of operators.
@@ 4. Statistical theory for regression of linear function-to-scalar maps
To still shed some light on this issue, we provide a detailed theoretical analysis of learning a linear PtO map $f$ that admits a factorization into a linear functional $q$, the QoI, composed with a self-adjoint linear operator $L$, the forward map. To set the stage, let $(H,\langle\cdot,\cdot\rangle,\lVert{\,\cdot\,}\rVert)$ be an infinite-dimensional real separable Hilbert space. We view the PtO map $f=q\circ L\colon H\to\mathbb{R}$ as a linear functional on $H$. Hence, the input space is always infinite-dimensional in this section. Let $\nu$ be the data-generating Borel probability measure on the input space $H$. Suppose that
@@
for some orthonormal basis $\{\varphi_{j}\}_{j\in\mathbb{N}}$ of $H$ and eigenvalue sequence $\{\sigma_{j}\}_{j\in\mathbb{N}}\subset\mathbb{R}_{\geq 0}$, and that $N$ iid input data samples $\{u_{n}\}_{n=1}^{N}\sim\nu^{\otimes N}$ are available. We consider two different supervised learning approaches (visualized in Figure 1 as F2V and F2F) that correspond to the given labeled output data $\{y_{n}\}_{n=1}^{N}$ being either noise-perturbed versions of
@@
In both cases, the primary goal is to estimate $f$ given input–output data pairs. Notice that in the (EE) approach, the responses are scalar-valued, while for the (FF) approach the responses are function-valued. [^2] Although the Hilbert space $H$ is general and not necessarily comprised of functions, we still refer to its elements as “functions” to avoid confusion caused by attempts to distinguish between finite-dimensional vectors and infinite-dimensional vectors. In some cases, the problem itself specifies the approach that can be used. For example, in physical experiments, it is often impossible to experimentally acquire the full output of $L$. Instead, only a finite number of possibly indirect measurements are available, which represent the QoI $q$ in the (EE) framework. Even if the data generation procedure is algorithmic, the cost of generating full-field data may be much higher than simply measuring finite-dimensional QoIs. Nevertheless, for a large class of forward maps $L$ and continuous QoIs $q$, we demonstrate that (FF) is more data efficient than (EE) in this linear setup; see insights (I1) and (I2) and Figure 2 for more details.
To this end, Subsection 4.1 sets up the framework for statistical error analysis of learning general linear functionals on $H$ with a particular Bayesian posterior estimator; this is the end-to-end setting. Subsection 4.2 then continues the setup for the setting of factorized linear functionals in the full-field learning setting. Subsection 4.3 contains the main results. Here, two theorems provide convergence rates for the end-to-end and full-field settings, respectively. In particular, the end-to-end result may be of independent interest to the functional data analysis and Bayesian nonparametric statistics communities. The theorems are followed by a discussion and a corollary that rigorously compares the data efficiency of end-to-end versus full-field learning of factorized linear functionals.
@@ 4.1. End-to-end learning
Consider a general linear functional $f\colon H\to\mathbb{R}$. Working in a nonparametric functional regression framework, we adopt a linear Gaussian posterior mean estimator that is obtained by conditioning a Gaussian process prior on the training dataset under the access model (EE). Our primary concern is the development of large sample convergence rates for the average squared prediction error of the estimator with respect to some test distribution $\nu^{\prime}$. We allow $\nu^{\prime}$ to be different from the input training distribution $\nu$; that is, our error bounds hold out-of-distribution or under covariate shift. We describe our statistical model and Bayesian inference approach in Subsection 4.1.1. In Subsection 4.1.2, we list and interpret the main assumptions that underlie the theory.
@@ 4.1.1. Setup and estimator
We adopt a Bayesian inverse problems perspective on the linear functional regression task. To simplify the analysis, suppose that $f\in H^{*}$ is continuous on $H$ [^3] It is possible to handle unbounded linear functionals $f$ using the weighted Hilbert–Schmidt approach from [16, Section 2.2] or the framework of measurable linear functionals from [29, Section 3 and 5], possibly at the expense of stronger assumptions on the data and prior covariances. As a mild abuse of notation, we use the same symbol $f$ for both the linear functional itself and its Riesz representer. We thus view $f$ as an element of $H$ in all that follows.
@@
$y_{n}=\langle f,u_{n}\rangle+\gamma\xi_{n}\,,\quad\text{where}\quad u_{n}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\nu\quad\text{and}\quad\xi_{n}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,1)\,.$ (12)
@@
Proceeding with the Bayesian approach, we endow $f$ with a Gaussian prior distribution $f\sim\mathcal{N}(0,\Lambda)$. Here $\Lambda\in\mathcal{L}(H)$ is a trace-class covariance operator on $H$. The advantage of working with the preceding linear Gaussian model is that the posterior distribution for $f$ is also a Gaussian measure supported on $H$ with closed form expressions for its posterior mean and covariance. To this end, let
$\widehat{\Sigma}\coloneqq\frac{S_{N}^{*}S_{N}}{N}=\frac{1}{N}\sum_{n=1}^{N}u_{n}\otimes u_{n}$ (14)
denote the empirical covariance operator of the input distribution $\nu$. Additionally, suppose that $U$, $\Xi$, and $f$ are independent as random variables. Although the randomness of the operator $S_{N}$ is a slight deviation from the usual Bayesian setting, application of [29, Proposition 3.1, pp. 2630–2631] and [15, Theorems 32, 13, and 37] still imply that the posterior distribution obtained by conditioning $f$ on the training dataset $(U,Y)$ is given by
@@
We take a frequentist consistency perspective by assuming there exists a fixed ground truth $f^{\dagger}\in H$ that generates the data $Y$ in (13). Abusing notation by using the same symbol for $Y$, this means that we observe the output response data [^4] The actual noise process $\Xi$ in the observed data (17) need not match the assumed Gaussian likelihood model implied by (12). Indeed, the main results in Subsection 4.3 remain valid for any centered square integrable random vector with isotropic covariance.
@@
Our estimator of $f^{\dagger}$ is then taken to be the posterior mean $\bar{f}^{(N)}=A_{N}Y$ from (15) with $Y$ as in (17). However, this particular estimator is chosen to simplify the exposition. As explained in Remark 4.9, the analysis remains valid if the full posterior distribution $\mathcal{N}(\bar{f}^{(N)},\Lambda^{(N)})$ from (15) is used to estimate $f^{\dagger}$ instead of just its mean. Posterior contraction rates also follow as a consequence.
Instead of measuring the quality of our estimate of $f^{\dagger}$ in the $H$-norm, we consider a weaker weighted norm induced by the average squared prediction error of the estimator. This is more common in statistical learning than in statistical inverse problems. To this end, let $\nu^{\prime}$ be a centered Borel probability measure supported on a sufficiently large space containing $H$. We are interested in the out-of-distribution test squared error of $\bar{f}^{(N)}$ with respect to $\nu^{\prime}$, which is given by
@@
Under the setting described in this section, the posterior mean is equivalent to a generalized-Tikhonov regularized least squares estimator. See [16, Section 2.3] and [29, pp. 3631–2632] for more details and relations to RKHS methods. This connects the Bayesian approach taken here back to traditional supervised learning frameworks.

### v2

@@ 1.4. Outline
The remainder of this article is organized as follows. We define the architecture of FNMs as a slight adjustment of FNOs in Section 2 (Contribution (C1)) and confirm that FNMs retain desirable properties of FNOs such as universal approximation in Section 3 (Contribution (C2)). In Section 4, we analyze end-to-end and full-field learning of linear functionals to establish a theoretical foundation that underlies the data volume requirements of the two approaches (Contribution (C3)); simulations also support the linear theory. Going beyond the theory, Section 5 provides numerical experiments that compare end-to-end and full-field learning with FNMs with both finite- and infinite-dimensional input space representations for predicting QoIs in several nonlinear PDE problems (Contribution (C4)). Concluding remarks are given in Section 6. Appendix A contains additional theorems related to Section 4. All proofs are provided in Appendices B and C.
2. Neural mappings for finite-dimensional vector data
@@ 2.1. A review of neural operators
Let $\mathcal{U}=\mathcal{U}(\mathcal{D};\mathbb{R}^{d_{u}})$ and $\mathcal{Y}=\mathcal{Y}(\mathcal{D};\mathbb{R}^{d_{y}})$ be Banach function spaces over Euclidean domain $\mathcal{D}\subset\mathbb{R}^{d}$. Finite-dimensional fully-connected neural networks are repeated compositions of affine mappings alternating with pointwise nonlinearities. To extend this framework to the infinite-dimensional function space setting, depth $T$ neural operators from $\mathcal{U}$ to $\mathcal{Y}$ take the form
@@
where $\mathcal{S}$ is a pointwise-defined local lifting operator, $\mathcal{Q}$ is a pointwise-defined local projection operator, and for each $t\in[T]$, the layer $\mathscr{L}_{t}\colon\mathcal{B}_{t-1}\to\mathcal{B}_{t}$ is a nonlinear map between appropriate Banach function spaces $\mathcal{B}_{t-1}(\mathcal{D};\mathbb{R}^{d_{t-1}})\subset L^{2}(\mathcal{D};\mathbb{R}^{d_{t-1}})$ and $\mathcal{B}_{t}(\mathcal{D};\mathbb{R}^{d_{t}})\subset L^{2}(\mathcal{D};\mathbb{R}^{d_{t}})$. The map $\mathscr{L}_{t}$ is the composition of a local (and usually nonlinear) operator with a nonlocal affine kernel integral operator [37].
The FNO is a specific instance of the class of neural operators (1). Let $\mathcal{D}=\mathbb{T}^{d}$. Then for FNO, the form of the layer $\mathscr{L}_{t}\colon\mathcal{B}_{t-1}(\mathbb{T}^{d};\mathbb{R}^{d_{t-1}})\to\mathcal{B}_{t}(\mathbb{T}^{d};\mathbb{R}^{d_{t}})$ is given by $v\mapsto\mathscr{L}_{t}(v)$, where for any $x\in\mathbb{T}^{d}$, it holds that
$\bigl{(}\mathscr{L}_{t}(v)\bigr{)}(x)=\sigma_{t}\bigl{(}W_{t}v(x)+(\mathcal{K}_{t}v)(x)+b_{t}(x)\bigr{)}\,.$ (2)
@@
In the preceding display, the $\psi_{k}=e^{2\pi\mathrm{i}\langle k,{\,\cdot\,}\rangle_{\mathbb{R}^{d}}}$ are the complex Fourier basis elements of $L^{2}(\mathbb{T}^{d};\mathbb{C})$ and $P^{(k)}_{t}\in\mathbb{C}^{d_{{t}}\times d_{t-1}}$ are the learnable parameters of the integral operator $\mathcal{K}_{t}$ for each $k\in\mathbb{Z}^{d}$. The functions $\sigma_{t}\colon\mathbb{R}\to\mathbb{R}$ are nonlinear activations that act pointwise when applied to vectors. Additional details of more general versions and computational implementations of the FNO may be found in [35, 37, 46].
Though the internal FNO layers $\{\mathscr{L}_{t}\}$ in (2) and (3) are defined on the periodic domain $\mathbb{T}^{d}$, it is possible to apply the FNO to other $d$-dimensional domains $\mathcal{D}\neq\mathbb{T}^{d}$. Define Banach spaces $\mathcal{B}_{\mathrm{in}}(\mathcal{D};\mathbb{R}^{d_{0}})$ and $\mathcal{B}_{\mathrm{out}}(\mathcal{D};\mathbb{R}^{d_{T}})$ and introduce an operator $\mathcal{E}\colon\mathcal{B}_{\mathrm{in}}\to\mathcal{B}_{0}$. Then, replace $\mathcal{S}$ in (1) with $\mathcal{E}\circ\mathcal{S}$. Similarly, let $\mathcal{R}\colon\mathcal{B}_{T}\to\mathcal{B}_{\mathrm{out}}$ be an operator that maps back to functions on the desired domain $\mathcal{D}$ and replace $\mathcal{Q}$ in (1) with $\mathcal{Q}\circ\mathcal{R}$. These modifications to the lifting and projecting components yield the final FNO architecture
@@ 2.2. The neural mappings framework
The neural operator architecture described in Section 2.1 only accepts inputs, outputs, and intermediate states that are elements of function spaces. Finite-dimensional vector inputs, outputs, and states are not directly compatible with neural operators. We propose neural mappings, which lift this restriction through two fundamental building blocks. The first, linear functional layers, map from function space to finite dimensions. The second, linear decoder layers, map from finite dimensions to function space. We combine these two building blocks with standard iterative neural operator layers to form several classes of nonlinear and function space consistent architectures.
Instating the neural operator notation from Section 2.1, we define a linear functional layer $\mathscr{G}\colon\mathcal{B}_{T-1}\to\mathbb{R}^{d_{T}}$ and a linear decoder layer $\mathscr{D}\colon\mathbb{R}^{d_{0}}\to\mathcal{B}_{1}$ to be maps of the form
@@
where we recall that $\{\psi_{k}\}$ is the Fourier basis of $L^{2}(\mathbb{T}^{d};\mathbb{C})$. In (6), the entries of the matrices $\{P^{(k)}\}\subset\mathbb{C}^{d_{T}\times d_{T-1}}$ correspond to the Fourier coefficients of the function $\kappa$ in (5). The calculation leading to the convergent series formula (6) uses Parseval’s theorem to equate the $L^{2}$ (5) and $\ell^{2}$ (6) inner products. Similar calculations show that, on the torus $\mathbb{T}^{d}$, the map $\mathscr{D}$ takes the form
@@
We are now in a position to define the general FNMs architecture.
@@ 3. Universal approximation theory for Fourier Neural Mappings
In this section, we establish universal approximation theorems for FNMs; this is a confirmation that the architectures maintain this desirable property of neural operators. The results are stated for the cases of the F2V and V2F architectures; the case of V2V trivially follows. Similar results also hold for general neural mappings by invoking the appropriate universal approximation theorems for general neural operators from [35, Section 9.3] and for the topology induced by Lebesgue–Bochner norms, i.e., average error with respect to a probability measure supported on the input space. For more details regarding these extensions, see [37, Theorems 11–14, Section 9.3, pp. 55–57] and [35, pp. 12–14 and Theorem 18]. Our proofs, which are collected in Appendix B, use arguments based on constant functions that are similar to those used to prove universal approximation theorems at the level of operators.
@@ 4. Statistical theory for regression of linear function-to-scalar maps
To still shed some light on this issue, we provide a detailed theoretical analysis of learning a linear PtO map $f$ that admits a factorization into a linear functional $q$, the QoI, composed with a self-adjoint linear operator $L$, the forward map. To set the stage, we adopt the following setup in the remainder of Section 4. Let $(H,\langle\cdot,\cdot\rangle,\lVert{\,\cdot\,}\rVert)$ be an infinite-dimensional real separable Hilbert space. We view the PtO map $f=q\circ L\colon H\to\mathbb{R}$ as a linear functional on $H$. Hence, the input space is always infinite-dimensional in this section. Let $\nu$ be the data-generating Borel probability measure on the input space $H$. We work in the setting that
@@
for some orthonormal basis $\{\varphi_{j}\}_{j\in\mathbb{N}}$ of $H$ and eigenvalue sequence $\{\sigma_{j}\}_{j\in\mathbb{N}}\subset\mathbb{R}_{\geq 0}$, and that $N$ i.i.d. input data samples $\{u_{n}\}_{n=1}^{N}\sim\nu^{\otimes N}$ are available.
We consider two different supervised learning approaches, visualized in Figure 1 as F2V and F2F, that correspond to the given labeled output training data $\{y_{n}\}_{n=1}^{N}$ being either noise-perturbed versions of
@@
In both cases, the primary goal is to estimate $f$ given input–output data pairs. Notice that in the (EE) approach, the responses are scalar-valued, while for the (FF) approach the responses are function-valued. [^2] Although the Hilbert space $H$ is general and not necessarily comprised of functions, we still refer to its elements as “functions” to avoid confusion caused by attempts to distinguish between finite-dimensional vectors and infinite-dimensional vectors. In some cases, the problem itself specifies the approach that can be used. For example, in physical experiments, it is often impossible to experimentally acquire the full output of $L$. Instead, only a finite number of possibly indirect measurements are available, which represent the QoI $q$ in the (EE) framework. Even if the data generation procedure is algorithmic, the cost of generating full-field data may be much higher than simply measuring finite-dimensional QoIs. Nevertheless, for a large class of forward maps $L$ and continuous QoIs $q$, we demonstrate that (FF) is more data-efficient than (EE) in this linear setup; see insights (I1) and (I2) and Figure 2 for more details.
To this end, Subsection 4.1 sets up the framework for statistical error analysis of learning general linear functionals on $H$ with a particular Bayesian posterior estimator; this is the end-to-end setting. Subsection 4.2 then continues the setup for the setting of factorized linear functionals in the full-field learning setting. Subsection 4.3 contains the main results. Here, two theorems provide convergence rates for the end-to-end and full-field settings, respectively. In particular, the end-to-end result may be of independent interest to the functional data analysis and Bayesian nonparametric statistics communities. The theorems are followed by a discussion and a corollary that directly compares the data efficiency of end-to-end versus full-field learning of factorized linear functionals. In Subsection 4.4, synthetic data experiments empirically confirm the theoretical results.
@@ 4.1. End-to-end learning
Consider a general linear functional $f\colon H\to\mathbb{R}$. Working in a nonparametric functional regression framework, we adopt a linear Gaussian posterior mean estimator that is obtained by conditioning a Gaussian process prior on the training dataset under the access model (EE). Our primary concern is the development of large sample convergence rates for the average squared prediction error of the estimator with respect to some centered test distribution $\nu^{\prime}$. We allow $\nu^{\prime}$ to be different from the input training distribution $\nu$ (11); that is, our error bounds hold out-of-distribution or under covariate shift. We describe our statistical model and Bayesian inference approach in Subsection 4.1.1. In Subsection 4.1.2, we list and interpret the main assumptions that underlie the theory.
@@ 4.1.1. Setup and estimator
We adopt a Bayesian inverse problems perspective on the linear functional regression task. To simplify the analysis, suppose that $f\in H^{*}$ is continuous on $H$ [^3] It is possible to handle unbounded linear functionals $f$ using the weighted Hilbert–Schmidt approach from [17, Section 2.2] or the framework of measurable linear functionals from [34, Section 3 and 5], possibly at the expense of stronger assumptions on the data and prior covariances. As a mild abuse of notation, we use the same symbol $f$ for both the linear functional itself and its Riesz representer. We thus view $f$ as an element of $H$ in all that follows.
@@
$y_{n}=\langle f,u_{n}\rangle+\gamma\xi_{n}\,,\quad\text{where}\quad u_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\nu\quad\text{and}\quad\xi_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)\,.$ (12)
@@
Proceeding with the Bayesian approach, we endow $f$ with a Gaussian prior distribution $f\sim\mathcal{N}(0,\Lambda)$. Here $\Lambda\in\mathcal{L}(H)$ is a trace-class covariance operator on $H$. The advantage of working with the preceding linear Gaussian model is that the posterior distribution for $f$ is also a Gaussian measure supported on $H$ with closed form expressions for its mean and covariance. To this end, let
$\widehat{\Sigma}\coloneqq\frac{S_{N}^{*}S_{N}^{\phantom{*}}}{N}=\frac{1}{N}\sum_{n=1}^{N}u_{n}\otimes u_{n}$ (14)
denote the empirical covariance operator of the centered input distribution $\nu$ (11). Additionally, suppose that $U$, $\Xi$, and $f$ are independent as random variables. Although the randomness of the operator $S_{N}$ is a slight deviation from the usual Bayesian setting, application of [34, Proposition 3.1, pp. 2630–2631] and [16, Theorems 32, 13, and 37] still imply that the posterior distribution obtained by conditioning the prior $f$ on the training dataset $(U,Y)$ is given by
@@
We take a frequentist consistency perspective by assuming there exists a fixed ground truth $f^{\dagger}\in H$ that generates the data $Y$ in (13). Abusing notation by using the same symbol for $Y$, this means that we observe the output response data [^4] The actual noise process $\Xi$ in the observed data (17) need not match the assumed Gaussian likelihood model implied by (12). Indeed, the main results in Subsection 4.3 for the posterior mean remain valid for any centered square integrable random vector with isotropic covariance.
@@
Our estimator of $f^{\dagger}$ is then taken to be the posterior mean $\bar{f}^{(N)}=A_{N}Y$ from (15) with $Y$ as in (17). However, the particular estimator $\bar{f}^{(N)}$ is chosen to simplify the exposition. As explained in Remark 4.10, the analysis remains valid if the full posterior distribution $\mathcal{N}(\bar{f}^{(N)},\Lambda^{(N)})$ from (15) is used to estimate $f^{\dagger}$ instead of just its mean. Posterior contraction rates also follow as a consequence.
Instead of measuring the quality of our estimate of $f^{\dagger}$ in the $H$-norm, we consider a weaker weighted norm induced by the average squared prediction error of the estimator. This weighted norm is more common in statistical learning than in statistical inverse problems; it includes the $H$-norm as a special case whenever $\operatorname{Cov}(\nu^{\prime})=\operatorname{Id}_{H}$. To this end, let $\nu^{\prime}$ be a centered Borel probability measure supported on a sufficiently large space containing $H$. We are interested in the out-of-distribution test squared error of $\bar{f}^{(N)}$ with respect to the test distribution $\nu^{\prime}$, which is given by
@@
Under the setting described in this section, the posterior mean is equivalent to a generalized-Tikhonov regularized least squares estimator. See [17, Section 2.3] and [34, pp. 3631–2632] for more details and relations to RKHS methods. This connects the Bayesian approach taken here back to traditional supervised learning frameworks.